OpenStudy (ellerai):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (0, –1) B. (0, 1) C. (7, –2) D. (3, –2)

OpenStudy (welshfella):

just replace the x and y in the equation by the values in each option to see which of them fit.

OpenStudy (ellerai):

would it be d?

OpenStudy (ellerai):

wait no

OpenStudy (ellerai):

c ?

OpenStudy (welshfella):

for example try the first choice 2x + 5y = 4 the pair (0,-1) means x = 0 and y = -1 so we have 2(0) + 5(-1) = -5 so its not this one

OpenStudy (ellerai):

okay thank you so much

OpenStudy (welshfella):

2(7) + 5(-2) = 14 - 10 = 4 yes

OpenStudy (welshfella):

C is correct